A boy rows his boat 8 km downstream in 1 hour and 4 minutes, and returns in 2 2/7 hours. At what rate will he row in still water? What is the rate of the current?
Answer by Alan3354(69443)

A boy rows his boat 8 km downstream in 1 hour and 4 minutes, and returns in 2 2/7 hours. At what rate will he row in still water? What is the rate of the current?
---------------
8 km/64 minutes = 1/8 km/min = 15/2 km/hr
8 km/(16/7) = 7/2 km/hr
---------------------
The boat's speed is the average, = (15/2 + 7/2)/2 = 5.5 km/hr
-------------
The current is the difference, = 2 km/hr